# A System Level Approach to LQR Control of the Diffusion Equation

### v1

@@ Abstract
The continuous-time, infinite horizon LQR problem for the diffusion equation over the unit circle with fully distributed actuation is considered. It is well-known that the solution to this problem can be obtained from the solution to an operator-valued algebraic Riccati equation. Here, it is demonstrated that this solution can be equivalently obtained by solving an $H_{2}$ control problem through a closed-loop design procedure that is analogous to the “System Level Synthesis” methodology previously developed for systems over a discrete spatial domain and/or over a finite time horizon. The presented extension to the continuous spatial domain and continuous and infinite-horizon time setting admits analytical solutions that may complement computational approaches for discrete or finite-horizon settings. It is further illustrated that spatio-temporal constraints on the closed-loop responses can be incorporated into this new formulation in a convex manner.
@@ I Introduction
As large-scale spatially-distributed systems with immense amounts of sensing and actuation become more prevalent, so does the need for tractable methods for optimal distributed controller design. In this setting, it is necessary to design control policies that utilize minimal communication between spatially distributed subcontroller units, each with access to a limited subset of system information. This design problem can be formulated as an optimal control problem subject to sparsity constraints that specify the measurement sharing and subcontroller communication structure of the system. It is well-known that such constrained optimization problems are non-convex except in special settings (quadratic invariance [1] or funnel causality [2]).
To overcome the computational complexity that arises from the non-convexity of these constrained optimization problems, various convex restrictions or relaxations have been developed, e.g., [3, 4]. One such convex approximation, the System Level Synthesis (SLS) framework [5], has gained significant recent attention. The SLS framework was originally formulated for finite-dimensional discrete-time systems over a finite time horizon (thus ensuring computational tractability) [5], but has since been adapted to various other settings including infinite time-horizons [6, Sec. 4], infinite spatial dimensions [7], and continuous-time dynamics [8]. Most recently, an SLS-like framework for distributed parameter systems over a continuous spatial domain was developed [9]; this result was restricted to finite-time horizons and discrete-time settings which led to tractable numerical solutions to the problem [9].
@@
The SLS framework is characterized by two key components: a synthesis step and an implementation step. The synthesis step relies on the fact that the cost functional of the design problem may be written as an affine function of well-chosen closed loop system mappings [10]. An “implementation” of the controller is then formed as an interconnection of elements that are built from these closed-loop maps, such that the controller inherits structural properties of the closed-loop mappings. (E.g., imposing that the closed-loops are FIR and sparse ensures that the resulting controller can also be implemented with sparse FIR elements.)
In this work, we demonstrate that an SLS-like procedure can recover the solution to the continuous-time, infinite-time-horizon Linear Quadratic Regulator (LQR) problem for a specific infinite-dimensional linear distributed parameter system: the diffusion equation. Although such optimal control problems have been extensively studied [11, 12], [13, Ch. 7], the formulation we present is novel.
@@
Analogous to the finite-dimensional SLS framework, we demonstrate that (1) the cost functional of interest can be written as an affine function of certain closed-loop mappings, and (2) the controller can be implemented by an interconnection of systems built from these closed-loop mappings. Although our results are presented for the diffusion equation, the approach presented may be modified to apply to other problem settings. This presents a first step toward developing a general SLS-like framework for distributed parameter systems in continuous-time and infinite time-horizon settings, which we believe will complement the computational results and theory developed in the discrete, finite-time horizon setting [9].
This paper is structured as follows. Section II provides notation and mathematical preliminaries. Section III introduces the diffusion dynamics and formulates the $H_{2}$ design problem of interest. Section IV presents our main result: an equivalent formulation of the proposed $H_{2}$ design problem in terms of the systems “closed-loop maps”, analogous to the SLS framework [5]. This problem is solved analytically in Section V, and the structure and stability properties of the resulting controller implementation are explored. The analytical solution is shown to recover the standard infinite-horizon LQR solution in Section VI. Finally, Section VII illustrates that constraints can be imposed on the closed-loop maps while preserving convexity of the underlying $H_{2}$ problem.
@@ II Notation & Mathematical Preliminaries
Let $\mathcal{X},\mathcal{U}$ denote two Hilbert spaces. $\mathcal{L}(\mathcal{X},\mathcal{U})$ denotes the set of linear operators from $\mathcal{X}$ to $\mathcal{U}$, and $\mathcal{L}(\mathcal{X})=\mathcal{L}(\mathcal{X},\mathcal{X}).$ An operator $B\in\mathcal{L}(\mathcal{X},\mathcal{U})$ is bounded if $\sup_{u\in\mathcal{U}}\frac{\|Bu\|_{\mathcal{X}}}{\|u\|_{\mathcal{U}}}<\infty$. $C^{\dagger}\in\mathcal{L}(\mathcal{U},\mathcal{X})$ denotes the adjoint of $C\in\mathcal{L}(\mathcal{X},\mathcal{U})$.
Let $\mathbb{T}$ denote the unit circle, $\mathbb{R}$ the real numbers, $\mathbb{C}$ the complex numbers, and $\mathbb{Z}$ the set of integers. Two Hilbert spaces of interest are the space $L_{n}^{2}(\mathbb{T})$ of $\mathbb{R}^{n}$-valued functions equipped with inner product $\left<f,g\right>_{L_{n}^{2}(\mathbb{T})}=\int_{\theta\in\mathbb{T}}f^{\top}(\theta)g(\theta)d\theta,$ and the space $\ell_{n}^{2}(\mathbb{Z})$ of $\mathbb{R}^{n}$-valued sequences equipped with inner product $\big<\hat{f},\hat{g}\big>_{\ell^{2}_{n}(\mathbb{Z})}=\sum_{\kappa\in\mathbb{Z}}\hat{f}^{\top}(\kappa)\hat{g}(\kappa).$ We neglect the subscript $n$ when clear from context.
Lower-case letters denote spatio-temporal signals over the domain $0,\infty)\times\mathbb{T},$ e.g.,
@@
Such signals can also be viewed as Hilbert-space valued signals over time, which we denote with bold-face lower case letters, e.g.,
${\boldsymbol{\psi}}(t):=\psi(t,\cdot).$ (2)
Let $\mathcal{X}$ be a Hilbert-space and $\boldsymbol{\psi}$}} an $\mathcal{X}$-valued signal. Define the signal ${{\boldsymbol{\psi}}}_{\beta}(t)=e^{-\beta t}{{\boldsymbol{\psi}}}(t).$ If ${{\boldsymbol{\psi}}}_{\beta}\in L_{1}([0,\infty),\mathcal{X})$, then $\boldsymbol{\psi}$}} is Laplace transformable with Laplace transform ${{\boldsymbol{\Psi}}}:\{s\in\mathbb{C};~{\rm Re}(s)\geq\beta\}\rightarrow\mathcal{X}$ given by [14, Ch. 7] ${\boldsymbol{\Psi}}(s):=\int_{0}^{\infty}e^{-st}{\boldsymbol{\psi}}(t)dt.$
$\mathcal{F}$ denotes the Fourier transform of elements of $L^{2}(\mathbb{T}):$
@@
which can be applied to the spatial component of spatio-temporal signals, e.g.,
$\hat{\psi}(t,j\kappa)=\frac{1}{2\pi}\int_{\theta\in\mathbb{T}}\psi(t,\theta)e^{-j\kappa\theta}d\theta.$ (4)
Recall that $\mathcal{F}$ is an isometric isomorphism to $\ell^{2}(\mathbb{Z})$.
A measurable function $\hat{A}:\mathbb{Z}\rightarrow\mathbb{C}^{m\times n}$, defines a multiplication operator $M_{\hat{A}}\in\mathcal{L}(\ell^{2}_{n}(\mathbb{Z}),\ell^{2}_{m}(\mathbb{Z}))$
$(M_{\hat{A}}\hat{f})(\kappa)=\hat{A}(\kappa)\hat{f}(\kappa).$ (5)
We refer to $\hat{A}_{\kappa}=\hat{A}(\kappa)$ as the symbol of the operator $M_{\hat{A}}$.
Letting ++ denote the addition operator on the group $\mathbb{T},$ for each $\xi\in\mathbb{T}$, define the translation operator $T_{\xi}\in\mathcal{L}(L^{2}_{n}(\mathbb{T})):$
$(T_{\xi}f)(\theta):=f(\theta+\xi).$ (6)
An operator $B\in\mathcal{L}(L^{2}_{m}(\mathbb{T}),L^{2}_{n}(\mathbb{T}))$ is translation invariant if it commutes with every translation operator, i.e. $T_{\xi}B=BT_{\xi}$ for all $\xi\in\mathbb{T}.$
@@ III Problem Set Up
$\tfrac{d}{dt}{{\boldsymbol{\psi}}}(t)$ $=(A{\boldsymbol{\psi}})(t)+(B{\bf u})(t)$ (7)
${\bf z}(t)$ $=(C{\boldsymbol{\psi}})(t)+(D{\bf u})(t),$
where $A$ generates a $C_{0}$-semigroup $\{T_{A}(t)\}$ on $\mathcal{D}(A)\subset L_{n}^{2}(\mathbb{T})$ and $B,C,D$ are elements of $\mathcal{L}(L^{2}(\mathbb{T}))$ spaces of compatible dimensions. We denote the system (7) by $\Sigma(A,B,C,D)$; we say that this system is strictly proper if $D\equiv 0$ and that this system is spatially invariant if each of $A,B,C,D$ is a translation invariant operator.
The semigroup $\{T_{A}(t)\}$ is exponentially stable if there exist constants $m,\eta>0$ for which
$\|T_{A}(t)\|_{L^{2}(\mathbb{T})\rightarrow L^{2}(\mathbb{T})}\leq me^{-\eta t},~~\text{for all}~t\geq 0.$ (8)
We define the infinite horizon observability Gramian, $L_{O}$, of the pair $(A,C)$ by [14]
$L_{O}=\int_{0}^{\infty}(T_{A}(t))^{\dagger}C^{\dagger}CT_{A}(t)dt,$ (9)
when this integral converges absolutely in operator norm. A sufficient condition for $L_{O}$ to be well-defined is that $C$ is bounded and that $\{T_{A}(t)\}$ is exponentially stable. We leverage $L_{O}$ to define the $H_{2}$ norm of a linear distributed parameter system.
@@
When $B$ is bounded and the infinite horizon observability Gramian $L_{O}$ is trace class, we define the $H_{2}$ norm of the system $\Sigma(A,B,C,D)$ as
$\|\Sigma(A,B,C,D)\|_{H_{2}}=\tfrac{1}{2\pi}{\rm Tr}(B^{\dagger}L_{O}B).$ (10)
Note that boundedness of $L_{O}$ is sufficient for this operator to be trace class; combined with the boundedness of $B$, this ensures the metric (10) will be finite.
We additionally utilize operator-valued transfer function representations of systems [14, Ch.7]. Let $\rho(A)$ denote the resolvent set and $\mathcal{D}(A)\subset L^{2}(\mathbb{T})$ denote the domain of the operator $A$. For $s\in\rho(A)$, the resolvent operator $R(s;A)=(sI-A)^{-1}$ is well-defined on $\mathcal{D}(A)$ by
$R(s;A)\psi_{0}=\int_{0}^{\infty}e^{-st}T_{A}(t)\psi_{0}dt,~\psi_{0}\in\mathcal{D}(A).$ (11)

### v2

@@ Abstract
The optimal controller design problem for a linear, first-order spatially-invariant distributed parameter system is considered. Through a case study of the Linear Quadratic Regulator (LQR) problem for the diffusion equation over the torus, it is illustrated that the optimal controller design problem can be equivalently formulated as an optimization problem over the system’s closed-loop mappings, analogous to the System Level Synthesis framework. This reformulation is solved analytically to recover the LQR for the diffusion equation, and an internally stable implementation of this controller is recovered from the optimal closed-loop mappings. It is further demonstrated that a class of spatio-temporal constraints on the closed-loop maps can be imposed on this closed-loop formulation while preserving convexity.
@@ I Introduction
As large-scale spatially-distributed systems with immense sensing and actuation become more prevalent, so does the need for tractable methods for optimal distributed controller design. One motivating example is the development of RF metamaterials for quantum computing, beamforming, and secure communications [23, 13, 21], which are modeled as continnum structures governed by PDEs. To leverage distributed sensing and actuation in these and other PDE systems, it is necessary to design control policies that use minimal communication between distributed subcontroller units, each with access to a limited subset of system information. This requires solving an optimal control problem subject to controller locality constraints, which is known to be non-convex except in special settings [18, 3].
@@
To overcome the computational complexity of such constrained optimization problems, various convex restrictions or relaxations have been developed, e.g., [19, 24]. Of particular significance is the System Level Synthesis (SLS) framework [22], which is characterized by (1) a synthesis step to formulate the design problem and constraints as affine in the system’s closed-loop mappings; and (2) an implementation step to construct an internally stable implementation of the controller from an interconnection of the optimal closed-loop maps. This implementation ensures that the controller inherits any structural properties constrained on the closed-loop maps. For example, imposing that the closed-loops are spatially sparse ensures that the resulting controller can also be implemented sparse elements.
The SLS framework was first formulated for finite-dimensional discrete-time systems over a finite time horizon [22], but has since been adapted to various other settings including infinite time-horizons [1, Sec. 4], countably infinite spatial dimensions [11], and continuous-time dynamics [12]. However, these approaches could not be directly applied to PDE systems formulated over a continuous spatial domain, as noted by the recent work of Conger et. al. [6]. Indeed, it is well-known that discretized approaches to design continuum controllers may lead to problematic “spillover effects” when the solutions to discretized formulations do not appropriately converge to the true continuum solution [16, 10] leading to excitement of problematic modes or even instability. This motivates the need for a development of an SLS framework in such continuous spatial domain settings.
The work of [6] presents a significant step in this direction. Notably, these results were restricted to finite-time horizons and discrete-time settings to enable tractable numerical solutions [6], but it remains unclear whether an analogous implementation of the controller can be obtained without such discrete-time and finite-time-horizon assumptions, while preserving internal stability. This work builds on the foundation in [6] to address this gap via an informative case study for control of the diffusion equation. The main results presented in this paper are as follows:
  • The infinite time horizon LQR problem for the diffusion equation over the torus is equivalently formulated as an optimization over the closed-loop mappings, analogous to the finite-dimensional SLS framework (Theorem 1).
  • This closed-loop formulation is solved analytically to recover the known LQR (Sections IV-C, VI).
  • An implementation of the optimal controller is constructed from the optimal closed-loop mappings, and is rigorously shown to preserve internal stability of the full infinite-dimensional system (Lemma 3).
  • A class of spatio-temporal constraints on the closed-loop mappings are shown to preserve convexity of the closed-loop design problem (Section VII).
@@
This work provides a step toward developing a general SLS-like framework for optimal controller design for distributed parameter systems in continuous-time and infinite time-horizon settings, which will complement the computational results and theory developed in [6]. Although the diffusion equation itself is not particularly interesting, it presents a proof of concept that can inform the use of similar methodology for other first-order in time linear, constant coefficient PDE systems.
@@ II Notation & Mathematical Preliminaries
Let $\mathbb{T}$ denote the unit circle, $\mathbb{R}$ the real numbers, $\mathbb{C}$ the complex numbers, and $\mathbb{Z}$ the integers. For $\mathcal{X},\mathcal{U}$ two Hilbert spaces, $\mathcal{L}(\mathcal{X},\mathcal{U})$ denotes the set of linear operators from $\mathcal{X}$ to $\mathcal{U}$, and $\mathcal{L}(\mathcal{X})=\mathcal{L}(\mathcal{X},\mathcal{X}).$ An operator $B\in\mathcal{L}(\mathcal{X},\mathcal{U})$ is bounded if $\sup_{u\in\mathcal{U}}\frac{\|Bu\|_{\mathcal{X}}}{\|u\|_{\mathcal{U}}}<\infty$. $C^{\dagger}\in\mathcal{L}(\mathcal{U},\mathcal{X})$ denotes the adjoint of $C\in\mathcal{L}(\mathcal{X},\mathcal{U})$. Two Hilbert spaces of interest are the space $L^{2}(\mathbb{T})$ equipped with inner product $\left<f,g\right>_{L^{2}(\mathbb{T})}=\int_{\theta\in\mathbb{T}}f^{*}(\theta)g(\theta)d\theta,$ and the space $\ell^{2}(\mathbb{Z})$ of sequences equipped with inner product $\big<\hat{f},\hat{g}\big>_{\ell^{2}(\mathbb{Z})}=\sum_{\kappa\in\mathbb{Z}}\hat{f}^{*}(\kappa)\hat{g}(\kappa).$
We consider spatio-temporal signals over the domain $0,\infty)\times\mathbb{T},$ which we denote by lower-case letters e.g.,
@@
We assume that such spatio-temporal signals $\psi(t,\cdot)\in L^{2}(\mathbb{T})$ for each $t$ and use bold-face lower case letters to denote the corresponding $L^{2}(\mathbb{T})$ valued signals, e.g., ${\boldsymbol{\psi}}(t):=\psi(t,\cdot).$
The Fourier transform, $\mathcal{F}$, of elements of $L^{2}(\mathbb{T}):$
@@
can be applied to the spatial component of the spatio-temporal signal $\psi(\cdot,\cdot)$:
$\hat{\psi}(t,j\kappa)=\frac{1}{2\pi}\int_{\theta\in\mathbb{T}}\psi(t,\theta)e^{-j\kappa\theta}d\theta,$ (3)
such that for each $t,$ $\boldsymbol{\psi}(t,\cdot)\in\ell^{2}(\mathbb{Z}).$
Letting ++ denote the addition operator on the group $\mathbb{T},$ for each $\xi\in\mathbb{T}$, define the translation operator $T_{\xi}\in\mathcal{L}(L^{2}(\mathbb{T})):$ $(T_{\xi}f)(\theta):=f(\theta+\xi).$ An operator $B\in\mathcal{L}(L^{2}(\mathbb{T}))$ is translation invariant if it commutes with every translation operator, i.e. $T_{\xi}B=BT_{\xi}$ for all $\xi\in\mathbb{T}.$ Such an operator can be represented in the Fourier domain as a multiplication operator, $B=\mathcal{F}^{-1}M_{\hat{B}}\mathcal{F}$ with $M_{\hat{B}}\in\mathcal{L}(\ell^{2}(\mathbb{Z}))$ such that
$(M_{\hat{B}}\hat{f})(\kappa)=\hat{B}(\kappa)\hat{f}(\kappa).$ (4)
We refer to the measurable function $\hat{B}_{\kappa}=\hat{B}(\kappa)$ as the symbol of the operator $M_{\hat{B}}$.
@@ III Problem Set Up
$\tfrac{d}{dt}{{\boldsymbol{\psi}}}(t)$ $=A{\boldsymbol{\psi}}(t)+B({\boldsymbol{u}}+{\boldsymbol{w}})(t)$ (5)
${\boldsymbol{z}}(t)$ $=C{\boldsymbol{\psi}}(t)+D{\boldsymbol{u}}(t),$
where $A,B,C,D$ are spatially-invariant operators, $A$ generates a $C_{0}$-semigroup $\{T_{A}(t)\}$ on $\mathcal{D}(A)\subset L^{2}(\mathbb{T})$, $B\in\mathcal{L}(L^{2}(\mathbb{T}))$, and $C,D\in\mathcal{L}(L^{2}(\mathbb{T}),L^{2}(\mathbb{T})\oplus L^{2}(\mathbb{T}))$. $\boldsymbol{u,w,\psi}$ and $\boldsymbol{z}$} denote the control signal, external disturbance, state, and performance output signal, respectively. We denote the system (5) by $\Sigma(A,B,C,D)$. A linear, spatially-invariant state feedback controller for this system takes the form
${\boldsymbol{u}}(t)=K\boldsymbol{\psi}(t),$ (6)
where $K\in\mathcal{L}(L^{2}(\mathbb{T}))$ is a translation invariant operator. We denote the system resulting from interconnection of (5) with controller $K$ as $\mathbb{F}(P;K)$; this system is also spatially invariant and is described by
$\tfrac{d}{dt}{{\boldsymbol{\psi}}}(t)$ $=(A+BK){\boldsymbol{\psi}}(t)+B{\bf w}(t)$ (7)
${\bf z}(t)$ $=(C+DK){\boldsymbol{\psi}}(t).$
The $C_{0}$-semigroup $\{T_{A_{\rm c.l.}}(t)\}$ generated by $A_{\rm c.l.}:=A+BK$ is exponentially stable if for some constants $m,\eta>0$,
$\|T_{A_{\rm c.l.}}(t)\|_{L^{2}(\mathbb{T})\rightarrow L^{2}(\mathbb{T})}\leq me^{-\eta t},~~\text{for all}~t\geq 0.$ (8)
@@
The infinite horizon observability Gramian, $L_{O}$, of the pair $(A_{\rm c.l.},C)$ is [7]
$L_{O}=\int_{0}^{\infty}(T_{A_{\rm c.l.}}(t))^{\dagger}C^{\dagger}CT_{A_{\rm c.l.}}(t)dt,$ (9)
when this integral converges absolutely in operator norm. A sufficient condition for $L_{O}$ to be well-defined is that $C$ is bounded and that $\{T_{A_{\rm c.l.}}(t)\}$ is exponentially stable. When $B$ is bounded and $L_{O}$ is trace class, the $\mathcal{H}_{2}$ norm [^1] Since we assumed ${\boldsymbol{w}}(t)\in L^{2}(\mathbb{T})$, we can equivalently interpret $\|\cdot\|_{\mathcal{H}_{2}}$ as the $L^{2}$ norm of the impulse response as outlined in [16, Ch. 5.1]. of the system $\mathbb{F}(P;K)=\Sigma(A_{\rm c.l.},B,(C+DK),0)$ is defined by
$\|\Sigma(A_{\rm c.l.},B,(C+DK),0)\|_{\mathcal{H}_{2}}=\tfrac{1}{2\pi}{\rm Tr}(B^{\dagger}L_{O}B).$ (10)
When $A_{\rm c.l.}$ generates an exponentially stable $C_{0}$-semigroup and $\|\Sigma(A_{\rm c.l.},B,(C+DK),0)\|_{\mathcal{H}_{2}}<\infty$, we say that the controller $K$ is internally stabilizing.
III-A Diffusion Dynamics
We focus on a specific plant - the diffusion equation over $\mathbb{T}$ with fully distributed actuation $u$ and disturbance $w$, described by the partial differential equation (PDE)
where $u(t,\cdot),w(t,\cdot),\psi(t,\cdot)\in L^{2}(\mathbb{T})$ and $\alpha>0$ is the constant of diffusivity. The spatio-temporal signal
${z}(t,\theta)=\left[\begin{array}[]{c}\gamma{\psi}(t,\theta)\\ {u}(t,\theta)\end{array}\right],~z(t,\cdot)\in L^{2}(\mathbb{T})\oplus L^{2}(\mathbb{T})$ (12)
is a performance output that provides a uniform penalty across the spatial domain with $\gamma>0$ determining a tradeoff between cost of state and cost of control. The interest in this system is not for its particular physical significance, but rather to serve as a case study for the proposed framework. Restricting to the torus is equivalent to enforcing periodic boundary conditions and will admit analytic solutions via Fourier analysis approaches [2].
The dynamics (11)-(12) can be written in the form (5) as
$\tfrac{d}{dt}{{\boldsymbol{\psi}}}(t)$ $=A{\boldsymbol{\psi}}(t)+{{\boldsymbol{u}}}(t)+{{\boldsymbol{w}}}(t)$ (13)
${{\boldsymbol{z}}}(t)$ $=C{\boldsymbol{\psi}}(t)+D{{\boldsymbol{u}}}(t),$
where $A\in\mathcal{L}(L^{2}(\mathbb{T}))$ is defined by $(A\psi)(t,\theta)=\alpha\partial_{\theta}^{2}\psi(t,\theta)$ and generates a $C_{0}$- semigroup $\{T_{A}(t)\}$ on
with derivatives, $f^{\prime},f^{\prime\prime}$, in (14) interpreted weakly, and $C,D\in\mathcal{L}(L^{2}(\mathbb{T}),L^{2}(\mathbb{T})\oplus L^{2}(\mathbb{T}))$ are the bounded operators
$(C\psi)(t,\theta)=\left[\begin{array}[]{c}\gamma\psi(t,\theta)\\ 0\end{array}\right],~(Du)(t,\theta)=\left[\begin{array}[]{c}0\\ u(t,\theta)\end{array}\right].$
$\Sigma(A,I,C,D)$ is a spatially-invariant system, equivalently characterized by its transformation under $\mathcal{F}$ as
$\tfrac{d}{dt}\hat{\psi}(t,\kappa)=-\alpha\kappa^{2}\hat{\psi}(t,\kappa)+\hat{u}(t,\kappa)+\hat{w}(t,\kappa),$ (15)
$\hat{z}(t,\kappa)=\underbrace{\left[\begin{array}[]{c}\gamma\\ 0\end{array}\right]}_{\hat{C}_{\kappa}}\hat{\psi}(t,\kappa)+\underbrace{\left[\begin{array}[]{c}0\\ 1\end{array}\right]}_{\hat{D}_{\kappa}}\hat{u}(t,\kappa),~t\in[0,\infty),~\kappa\in\mathbb{Z}.$
Note that for each fixed $\kappa\in\mathbb{Z},$ (LABEL:eq:open_loop_kappa) is a system with finitely many inputs and outputs.
III-B Optimal Control Formulations
We consider the Linear Quadratic Regulator (LQR) problem for the diffusion equation (13) [4, 7]:
$\inf_{u}~\int_{0}^{\infty}\left<{\boldsymbol{\psi}}(t),\gamma^{2}{\boldsymbol{\psi}}(t)\right>_{L^{2}(\mathbb{T})}+\left<{{\boldsymbol{u}}}(t),{{\boldsymbol{u}}}(t)\right>_{L^{2}(\mathbb{T})}dt$ (16)
$~{\rm s.t.}~~\partial_{t}\psi(t,\theta)=\alpha\partial_{\theta}^{2}\psi(t,\theta)+u(t,\theta),~{{\boldsymbol{\psi}}}(0)=\psi_{0}.$
The solution $K^{\rm opt}$ to (LABEL:eq:LQR) is independent of initial condition $\psi_{0}\in L^{2}(\mathbb{T})$ and it will be shown in Section VI that $K^{\rm opt}$ is the solution to the following $\mathcal{H}_{2}$ problem:
$K^{\rm opt}=$ $\operatorname*{arg\,min}_{K}~~\|\mathbb{F}(P;K)\|_{\mathcal{H}_{2}}$ (17)
$~~~~{\rm s.t.}~~~~K~\text{spatially-invariant}$
$~~~~~~~~~~~~K~\text{internally stabilizing}.$
Remark 1
In some cases (LABEL:eq:LQR) and (17) result in the same $K^{\rm opt}$, but this does not always hold for PDE systems. E.g., if operation by $K$ requires convolution with a delta distribution, the impulse response of the closed-loop system will not be $L^{2}(\mathbb{T})$; thus, the $\mathcal{H}_{2}$ problem is not well-posed, although the LQR problem is well-posed with finite cost $\forall~\psi_{0}\in\mathcal{D}(A)$.
IV Closed-Loop Formulation of $\mathcal{H}_{2}$ Problem
In this section, we demonstrate that the controller design problem (17) can be formulated as an optimization problem over certain closed-loop mappings and admits analytical solutions in the spatial frequency domain. We will illustrate that this formulation allows for spatio-temporal constraints to be imposed in a convex manner in Section VII.
IV-A Transfer Functions
To introduce closed-loop mappings, we use operator-valued transfer function representations of systems [7, Ch.7] to denote mappings between temporally Laplace transformed variables. Under certain technical conditions [7, Ch. 7], for some positive constant $\beta$, the Laplace transform of a signal $\boldsymbol{\psi}$}}, ${{\boldsymbol{\Psi}}}:\{s\in\mathbb{C};~{\rm Re}(s)\geq\beta\}\rightarrow L^{2}(\mathbb{T})$, is well-defined by ${\boldsymbol{\Psi}}(s):=\int_{0}^{\infty}e^{-st}{\boldsymbol{\psi}}(t)dt.$ First let $\Sigma(A,B,C,D)$ be a general system (5). Let $\rho(A)$ denote the resolvent set and $\mathcal{D}(A)\subset L^{2}(\mathbb{T})$ denote the domain of the operator $A$. For $s\in\rho(A)$, the resolvent operator $(sI-A)^{-1}$ is well-defined on $\mathcal{D}(A)$ by
$(sI-A)^{-1}\psi_{0}:=\int_{0}^{\infty}e^{-st}T_{A}(t)\psi_{0}dt,~\psi_{0}\in\mathcal{D}(A).$ (18)